OpenStudy (nesha97):

16. If you deposit $5,000 in an account that pays 5% interest compounded annually, how much money will you have in your account at the end of 15 years? Write an exponential function that represents this situation.

OpenStudy (nesha97):

if i did it right i came out with 5263.16 (rounded)

jimthompson5910 (jim_thompson5910):

Did you use this formula? \[\LARGE A = P*\left(1+\frac{r}{n}\right)^{n*t}\]

OpenStudy (nesha97):

so p would be the percent a is what ever number i get n is the years???? and r is the first amount (5000)???? and t idk what t would be

jimthompson5910 (jim_thompson5910):

P = principal = amount deposited = 5000 r = interest rate in decimal form = 0.05 n = number of times you compound a year = 1 (annual compounding = 1 time a year) t = 15 years

jimthompson5910 (jim_thompson5910):

In summary of the values you plug in P = 5000 r = 0.05 n = 1 t = 15

OpenStudy (nesha97):

i got 10400 when i did that

OpenStudy (nesha97):

thank you so much

OpenStudy (nesha97):

can you help me out with this one also \[a ^{m}\times a ^{3}= \frac{ 1 }{ a ^{2} }\] it ask me what a is and im confuse because a can be anything right?

jimthompson5910 (jim_thompson5910):

are you sure it's not asking about m? Or is it asking about restrictions on 'a'?

OpenStudy (nesha97):

yes im sure, I thought it did at first but when i reread it was a

jimthompson5910 (jim_thompson5910):

well the exponents m and 3 add up to m+3 so we have a^(m+3) = 1/(a^2) a^(m+3) = a^(-2) which means m+3 = -2 leading to m = -5 If m = -5, then 'a' can be any number you want as long as 'a' is not zero

jimthompson5910 (jim_thompson5910):

I'm not sure what your teacher is looking for to be honest

OpenStudy (nesha97):

yea i know I can email her in the morning and ask if she meant m and not a

OpenStudy (nesha97):

thank you

jimthompson5910 (jim_thompson5910):

no problem